# Superlinear Convergence of an Interior Point Algorithm on Linear Semi-definite Feasibility Problems with Application to Linear Matrix Inequalities

### v3

@@ 3 Homogeneous Feasibility Model as a Semi-definite Linear Complementarity Problem
Note that the above SDLCP representation of the homogeneous feasibility model has the structure of an LSDFP. Recall that an LSDFP written as the SDLCP (20)-(22) is such that in (20), $(\mathcal{A}_{1}(X^{1}))_{i}=0$ for $i=m_{1}+1,\ldots,\tilde{n}_{1}$, $(\mathcal{B}_{1}(Y^{1}))_{j}=0$ for $j=1,\ldots,m_{1}$, and $q_{i}=0$ for $i=1,\ldots,m_{1}$ or $q_{i}=0$ for $i=m_{1}+1,\ldots,\tilde{n}_{1}$. Here, $q=(q_{1},\ldots,q_{\tilde{n}_{1}})^{T}$. From (80), (85), we see that for our SDLCP representation, it has this structure with $m_{1}=m+n$ and $q=0$. Having $q=0$, in particular, $q_{i}=0$ for $i=1,\ldots,m_{1}$, is the property satisfied by our SDLCP representation that does not hold for the LSDFP considered in this paper. In fact, having this property is the key that allows us to apply results in the literature, namely [26], to show the main result (Theorem 4.1) in this paper on superlinear convergence.
The following proposition relates a solution of the homogeneous feasibility model of the LSDFP to that of its SDLCP representation:
@@
$X^{1}=\left[\begin{array}[]{cc}X&0\\ 0&\tau\end{array}\right],\quad Y^{1}=\left[\begin{array}[]{cc}Y&\ast\\ \ast&\kappa\end{array}\right],$ (90)
where the “normal-$\ast$” entries in $Y^{1}$ are such that $Y^{1}\in S^{n_{1}}_{+}$, satisfies (20)-(22), with $n_{1}=n+1$, $q=0$, $\mathcal{A}_{1},\mathcal{B}_{1}$ are given by (80), (85) respectively. On the other hand, if $(X^{1},Y^{1})$ satisfies (20)-(22), where $n_{1}=n+1$, $q=0$, $\mathcal{A}_{1},\mathcal{B}_{1}$ are given by (80), (85) respectively, then $X^{1},Y^{1}$ are given by (90), and $(X,Y,\tau,\kappa)$ satisfies (74), (67).
Proof: The proposition is clear based on how $\mathcal{A}_{1}$ and $\mathcal{B}_{1}$ are defined in (80) and (85) respectively.
We have another proposition below that shows that if the LSDFP satisfies Assumption 2.1, then the SDLCP representation of its homogeneous feasibility model satisfies Assumption 3.1:
@@
The SDLCP (20)-(22) with $n_{1}=n+1$, $q=0$ and $\mathcal{A}_{1}$, $\mathcal{B}_{1}$ given by (80), (85) respectively, satisfies Assumption 3.1 when the LSDFP satisfies Assumption 2.1.
@@
Given the SDLCP (20)-(22) with $n_{1}=n+1$, $q=0$ and $\mathcal{A}_{1}$, $\mathcal{B}_{1}$ given by (80), (85) respectively. Suppose
${\mathcal{A}}_{1}(X^{1})+{\mathcal{B}}_{1}(Y^{1})=0$
for some $(X^{1},Y^{1})\in S^{n_{1}}\times S^{n_{1}}$. Then $X^{1}$ and $Y^{1}$ are given by
$X^{1}=\left[\begin{array}[]{cc}X&0\\ 0&\tau\end{array}\right],\quad Y^{1}=\left[\begin{array}[]{cc}Y&\ast\\ \ast&\kappa\end{array}\right],$
where $(X,Y,\tau,\kappa)$ satisfies (74). Hence, by the first observation above, we have ${\rm{Tr}}(X^{1}Y^{1})={\rm{Tr}}(XY)+\tau\kappa=0$. Therefore, Assumption 3.1(a) holds. Furthermore, Assumption 3.1(b) holds since a solution to the given SDLCP is $X^{1}=0,Y^{1}=0$. Finally, the second observation above means that the matrix
@@
has full row rank. This implies that the matrix $(\mathcal{A}_{1}\ \mathcal{B}_{1})$, where $\mathcal{A}_{1}$ and $\mathcal{B}_{1}$ are defined by (80) and (85) respectively, has full row rank, and hence Assumption 3.1(c) holds as well.
@@ 4 An Infeasible Interior Point Algorithm on the Homogeneous Feasibility Model
We describe in this section an infeasible path-following interior point algorithm, Algorithm 4.1, on the homogeneous feasibility model (16)-(19) (or (66), (67) or (74), (67)). It generates iterates that follow an infeasible central path in a (narrow) neighborhood. This algorithm is a predictor-corrector type algorithm on the homogeneous feasibility model, and is first considered in [21].
@@
In the algorithm, Algorithm 4.1, which is described below, iterates generated by the algorithm always stay within a neighborhood of the central path. We consider the dual Helmberg-Kojima-Monteiro (HKM) search direction in the algorithm, even though our results also hold for the Nesterov-Todd (NT) search direction - see Remark 4.3. Among different search directions used in interior point algorithms on SDLCPs/SDPs, the Alizadeh-Haeberly-Overton (AHO) [1], Helmberg-Kojima-Monteiro (HKM) [9, 12, 15] and Nesterov-Todd (NT) [17, 18] search directions are better known, with the latter two being implemented in SDP solvers, such as SeDuMi [29] and SDPT3 [30]. It is worth noting that the HKM direction has also been implemented in the SDPA package to solve SDPs [31, 32].
@@
(See Algorithm 5.1 of [21]) Given $\beta_{1}<\beta_{2}$ with $\beta_{2}^{2}/(2(1-\beta_{2})^{2})\leq\beta_{1}<\beta_{2}<\beta_{2}/(1-\beta_{2})<1$. Choose $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ with $(n+1)\mu_{0}={\rm{Tr}}(X_{0}Y_{0})+\tau_{0}\kappa_{0}$. For $k=0,1,\ldots$, perform ($a1$) through ($a5$):
@@
If $\max\{({\rm{Tr}}(XY)+\tau\kappa)/\tau^{2},|r_{1}/\tau|,\ldots,|r_{m}/\tau|,\|s/\tau\|\}\leq\epsilon$, then report $(X/\tau,y/\tau,Y/\tau)$ as an approximate solution to the LSDFP, and terminate. If $\tau$ is sufficiently small, terminate with no optimal solutions to the LSDFP.
@@
[Corrector Step] Find the solution $(\Delta X_{c},\Delta y_{c},\Delta Y_{c},\Delta\tau_{c},\Delta\kappa_{c})$ of the linear system (93)-(97), with $\sigma=1-\overline{\alpha}$, $\overline{r}_{i}=0,i=1,\ldots,m$, $\overline{s}=0$ and $\overline{\gamma}=0$. Set
@@
Polynomial iteration complexity for a general version of Algorithm 4.1 used to solve (11)-(14) is shown in [21] where the HKM direction is considered. The result there also holds for the dual HKM direction. Hence, Algorithm 4.1 has polynomial iteration complexity when it is used to solve (16)-(19). Furthermore, we remark that Algorithm 4.1 can easily be adapted to solve the LSDFP instead of its homogeneous feasibility model. An advantage of applying the algorithm on its homogeneous feasibility model is that we have guaranteed superlinear convergence of iterates generated by the algorithm, as shown in Theorem 4.1.
For $k=0,1,\ldots$, define
$(r_{k})_{i}$ ::=:= ${\rm{Tr}}(A_{i}X_{k})-b_{i}\tau_{k},\quad i=1,\ldots,m,$
$s_{k}$ ::=:= $\sum_{i=1}^{m}(y_{k})_{i}A_{i}+Y_{k},$
$\gamma_{k}$ ::=:= $\kappa_{k}-b^{T}y_{k}.$
@@
Throughout this paper, we consider Algorithm 4.1 with initial iterate $(X_{0},y_{0},Y_{0},$ $\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ such that $(y_{0},Y_{0})\in\Re^{m}\times S^{n}_{++}$ is feasible to the dual SDP (8) with $C=0$. Therefore, in the algorithm, we have $s_{0}=0$ while $(r_{0})_{i},i=1,\ldots,m$, and $\gamma_{0}$ are generally nonzero.
Algorithm 4.1 can be written in an equivalent way as Algorithm 4.2, which we present below. We use Algorithm 4.2 to solve the SDLCP representation of the homogeneous feasibility model, (20)-(22), where $n_{1}=n+1$, $q=0$ and $\mathcal{A}_{1},\mathcal{B}_{1}$ are given by (80), (85) respectively. Before describing the algorithm, we first define an analogous (narrow) neighborhood of the central path of the SDLCP representation:
$\mathcal{N}_{1}(\beta,\mu^{1}):=\{(X^{1},Y^{1})\in S^{n_{1}}_{++}\times S^{n_{1}}_{++}\ ;\ \|(Y^{1})^{1/2}X^{1}(Y^{1})^{1/2}-\mu^{1}I\|_{F}\leq\beta\mu^{1},\mu^{1}={\rm{Tr}}(X^{1}Y^{1})/n_{1}\}.$
We again have a system of equations for $(\Delta X^{1},\Delta Y^{1})\in S^{n_{1}}\times S^{n_{1}}$ that plays an important role in the algorithm, just like the system of equations (93)-(97) for Algorithm 4.1:
$(Y^{1})^{1/2}(X^{1}\Delta Y^{1}+\Delta X^{1}Y^{1})(Y^{1})^{-1/2}+(Y^{1})^{-1/2}(\Delta Y^{1}X^{1}+Y^{1}\Delta X^{1})(Y^{1})^{1/2}$
$=2(\sigma\mu^{1}I-(Y^{1})^{1/2}X^{1}(Y^{1})^{1/2}),$ (111)
$\mathcal{A}_{1}(\Delta X^{1})+\mathcal{B}_{1}(\Delta Y^{1})=-\overline{r}.$ (112)
Note that in Algorithm 4.2, it can be shown that $(\Delta X^{1},\Delta Y^{1})$ obtained by solving (111), (112) always exists and is unique [23, 26].
@@
(See Algorithm 4.1 of [26]; Algorithm 2.1 of [23]) Given $\beta_{1}<\beta_{2}$ with $\beta_{2}^{2}/(2(1-\beta_{2})^{2})\leq\beta_{1}<\beta_{2}<\beta_{2}/(1-\beta_{2})<1$. Choose $(X_{0}^{1},Y_{0}^{1})\in S^{n_{1}}_{++}\times S^{n_{1}}_{++}$ such that
$X^{1}_{0}=\left[\begin{array}[]{cc}X_{0}&0\\ 0&\tau_{0}\end{array}\right],\quad Y^{1}_{0}=\left[\begin{array}[]{cc}Y_{0}&0\\ 0&\kappa_{0}\end{array}\right],$ (117)
@@
Set $X^{1}=X^{1}_{k}$, $Y^{1}=Y^{1}_{k}$, and define
$r:=\mathcal{A}_{1}(X^{1})+\mathcal{B}_{1}(Y^{1}).$
@@
If $\max\{{\rm{Tr}}(X^{1}Y^{1})/(X^{1})^{2}_{n_{1},n_{1}},\|r/X^{1}_{n_{1},n_{1}}\|\}\leq\epsilon$, then terminate with the solution $(X^{1},Y^{1})$. If $X^{1}_{n_{1},n_{1}}$ is sufficiently small, terminate with no optimal solutions to the LSDFP.
@@
[Predictor Step] Find the solution $(\Delta X^{1}_{p},\Delta Y^{1}_{p})$ of the linear system (111), (112), with $\sigma=0$, $\overline{r}=r$.
@@
$\overline{X}^{1}=X^{1}+\overline{\alpha}\Delta X^{1}_{p},\ \ \overline{Y}^{1}=Y^{1}+\overline{\alpha}\Delta Y^{1}_{p},$
@@
$\delta$ $=$ $\frac{1}{\mu^{1}}\|(Y^{1})^{1/2}\Delta X_{p}^{1}\Delta Y_{p}^{1}(Y^{1})^{-1/2}\|_{F},$ (120)
@@
$\alpha_{2}=\max\{\tilde{\alpha}\in[0,1]\ ;\ (X^{1}+\alpha\Delta X^{1}_{p},Y^{1}+\alpha\Delta Y^{1}_{p})\in\mathcal{N}_{1}(\beta_{2},(1-\alpha)\mu^{1})\ \forall\ \alpha\in[0,\tilde{\alpha}]\}.$

### v4

@@ 3 Homogeneous Feasibility Model as a Semi-definite Linear Complementarity Problem
Note that the above SDLCP representation of the homogeneous feasibility model has the structure of an LSDFP. Recall that an LSDFP written as the SDLCP (20)-(22) is such that in (20), $({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}}))_{i}=0$ for $i=m_{1}+1,\ldots,\tilde{n}_{1}$, (^$({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}}))_{j}=0$ for $j=1,\ldots,m_{1}$, and $q_{i}=0$ for $i=1,\ldots,m_{1}$ or $q_{i}=0$ for $i=m_{1}+1,\ldots,\tilde{n}_{1}$. Here, $q=(q_{1},\ldots,q_{\tilde{n}_{1}})^{T}$. From (73), (78), we see that for our SDLCP representation, it has this structure with $m_{1}=m+n$ and $q=0$. Having $q=0$, in particular, $q_{i}=0$ for $i=1,\ldots,m_{1}$, is the property satisfied by our SDLCP representation, sSDLCP, that does not hold for the sLSDFP. In fact, having this property is the key that allows us to apply results in the literature, namely [26], to show the main result (Theorem 4.1) in this paper on superlinear convergence.
The following proposition relates a solution of the homogeneous feasibility model of the sLSDFP to that of its SDLCP representation, sSDLCP:
@@
${{\color[rgb]{0,0,0}\hat{X}}}=\left[\begin{array}[]{cc}X&0\\ 0&\tau\end{array}\right],\quad{{\color[rgb]{0,0,0}\hat{Y}}}=\left[\begin{array}[]{cc}Y&\ast\\ \ast&\kappa\end{array}\right],$ (83)
where the $\ast$” entries in ${{\color[rgb]{0,0,0}\hat{Y}}}$ are such that ${{\color[rgb]{0,0,0}\hat{Y}}}\in S^{n_{1}}_{+}$, satisfies sSDLCP. On the other hand, if $({{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}})$ satisfies sSDLCP, then ${{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}}$ are given by (83), and $(X,Y,\tau,\kappa)$ satisfies (66), (67).
Proof: The proposition is clear based on how ${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}$ and $^{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}$ are defined in (73) and (78) respectively.
The following assumptions are imposed on the SDLCP (20)-(22), which we show in Proposition 3.2 to hold for the SDLCP representation, sSDLCP, of the homogeneous feasibility model of the sLSDFP.
Assumption 3.1
  1. (a)
    System (20)-(22) is monotone. That is, ${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}})=0$ for ${{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}}\in S^{n_{1}}\Rightarrow{\rm{Tr}}({{\color[rgb]{0,0,0}\hat{X}}}{{\color[rgb]{0,0,0}\hat{Y}}})\geq 0$.
  2. (b)
    There exists at least one solution to SDLCP (20)-(22).
  3. (c)
    $\{{{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}})\ ;\ {{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}}\in S^{n_{1}}\}=\Re^{\tilde{n}_{1}}$.
Assumption 3.2
There exist ${{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}}\in S^{n_{1}}_{++}$ such that ${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}})=q$.
Assumptions 3.1(a), (c) and 3.2 are imposed in [12] where the paper studies feasible interior point algorithms on the SDLCP (20)-(22), while [10, 26] assume Assumption 3.1 in the study of infeasible interior point algorithms on the SDLCP (20)-(22). In the study of infeasible interior point algorithms on the SDLCP (20)-(22), it is not necessary to impose Assumption 3.2. A reason is that we do not need a strictly feasible initial iterate to the algorithm. To solve the primal-dual SDP pair (4)-(8) by applying an interior point algorithm on its homogeneous feasibility model (11)-(14), the algorithm is necessarily infeasible as discussed near the end of Section 2. When we express the homogeneous feasibility model as an SDLCP, the interior point algorithm that is applied to the homogeneous feasibility model can be equivalently applied to the corresponding SDLCP, and by necessity, the algorithm on the SDLCP is infeasible just like that for the homogeneous feasibility model. Therefore, having Assumption 3.2 imposed on the SDLCP is not suitable and in fact can never hold in our case as there cannot exist an $(X,Y)\in S^{n}_{++}\times S^{n}_{++}$ such that (20) holds for the SDLCP obtained from the homogeneous feasibility model. We therefore have Assumption 3.1(b) in its place. In the context when the primal-dual SDP pair (4)-(8) has $C=0$, we see in Proposition 3.2 below that when the sLSDFP satisfies Assumption 2.1, the SDLCP representation, sSDLCP, of the homogeneous feasibility model (16)-(19) satisfies Assumption 3.1.
@@
We have the sSDLCP satisfies Assumption 3.1 when the the sLSDFP satisfies Assumption 2.1.
@@
Given the sSDLCP. Suppose
${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}})=0$
for some $({{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}})\in S^{n_{1}}\times S^{n_{1}}$. Then ${{\color[rgb]{0,0,0}\hat{X}}}$ and ${{\color[rgb]{0,0,0}\hat{Y}}}$ are given by
${{\color[rgb]{0,0,0}\hat{X}}}=\left[\begin{array}[]{cc}X&0\\ 0&\tau\end{array}\right],\quad{{\color[rgb]{0,0,0}\hat{Y}}}=\left[\begin{array}[]{cc}Y&\ast\\ \ast&\kappa\end{array}\right],$
where $(X,Y,\tau,\kappa)$ satisfies (66). Hence, by the first observation above, we have ${\rm{Tr}}({{\color[rgb]{0,0,0}\hat{X}}}{{\color[rgb]{0,0,0}\hat{Y}}})={\rm{Tr}}(XY)+\tau\kappa=0$. Therefore, Assumption 3.1(a) holds, since $\hat{A}(\hat{X})+\hat{B}(\hat{Y})=0$ for $\hat{X},\hat{Y}\in S^{n_{1}}$ implies that ${\mbox{Tr}}(\hat{X}\hat{Y})=0$, as argued above. Furthermore, Assumption 3.1(b) holds since a solution to the given sSDLCP is ${{\color[rgb]{0,0,0}\hat{X}}}=0,{{\color[rgb]{0,0,0}\hat{Y}}}=0$. Finally, the second observation above means that the matrix
@@
has full row rank. This implies that the matrix $({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}\ {{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})$, where ${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}$ and $^{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}$ are defined by (73) and (78) respectively, has full row rank, and hence Assumption 3.1(c) holds as well.
@@ 4 An Infeasible Interior Point Algorithm on the Homogeneous Feasibility Model
We describe in this section an infeasible path-following interior point algorithm, Algorithm 4.1, on the homogeneous feasibility model (16)-(19) (or (66), (67)). It generates iterates that follow an infeasible central path in a (narrow) neighborhood. This algorithm is a predictor-corrector type algorithm on the homogeneous feasibility model, and is first considered in [21].
@@
In Algorithm 4.1, which is described below, iterates always stay within a neighborhood of the central path. We consider the dual Helmberg-Kojima-Monteiro (HKM) search direction in the algorithm, even though our results also hold for the Nesterov-Todd (NT) search direction - see Remark 4.3. Among different search directions used in interior point algorithms on SDLCPs/SDPs, the Alizadeh-Haeberly-Overton (AHO) [1], Helmberg-Kojima-Monteiro (HKM) [9, 12, 15] and Nesterov-Todd (NT) [17, 18] search directions are better known, with the latter two being implemented in SDP solvers, such as SeDuMi [29] and SDPT3 [30]. It is worth noting that the HKM direction has also been implemented in the SDPA package to solve SDPs [31, 32].
@@
(See Algorithm 5.1 of [21]) Given $\epsilon>0$, and $\beta_{1}<\beta_{2}$ with $\beta_{2}^{2}/(2(1-\beta_{2})^{2})\leq\beta_{1}<\beta_{2}<\beta_{2}/(1-\beta_{2})<1$. Choose $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ with $(n+1)\mu_{0}={\rm{Tr}}(X_{0}Y_{0})+\tau_{0}\kappa_{0}$. For $k=0,1,\ldots$, perform ($a1$) through ($a5$):
@@
If $\max\{({\rm{Tr}}(XY)+\tau\kappa)/\tau^{2},|r_{1}/\tau|,\ldots,|r_{m}/\tau|,\|s/\tau\|\}\leq\epsilon$, then report $(X/\tau,y/\tau,Y/\tau)$ as an approximate solution to the sLSDFP, and terminate. If $\tau$ is sufficiently small, terminate with no optimal solutions to the sLSDFP.
@@
[Corrector Step] Find the solution $(\Delta X_{c},\Delta y_{c},\Delta Y_{c},\Delta\tau_{c},\Delta\kappa_{c})$ of the linear system (86)-(90), with ${{\color[rgb]{0,0,0}X=\overline{X},y=\overline{y},Y=\overline{Y},\tau=\overline{\tau},\kappa=\overline{\kappa}}},\sigma=1-\overline{\alpha}$, $\overline{r}_{i}=0,i=1,\ldots,m$, $\overline{s}=0$ and $\overline{\gamma}=0$. Set
@@
Polynomial iteration complexity for a general version of Algorithm 4.1 used to solve (11)-(14) is shown in [21] where the HKM direction is considered. The result there also holds for the dual HKM direction. Hence, Algorithm 4.1 has polynomial iteration complexity when it is used to solve (16)-(19). Furthermore, we remark that Algorithm 4.1 can easily be adapted to solve the sLSDFP instead of its homogeneous feasibility model. An advantage of applying the algorithm to its homogeneous feasibility model is that we have guaranteed superlinear convergence of iterates generated by the algorithm, as shown in Theorem 4.1.
@@
Throughout this paper, we consider Algorithm 4.1 with initial iterate $(X_{0},y_{0},Y_{0},$ $\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ such that $(y_{0},Y_{0})\in\Re^{m}\times S^{n}_{++}$ is feasible to the sDSDP. Therefore, in the algorithm, we have $\sum_{i=1}^{m}(y_{0})_{i}A_{i}+Y_{0}=0$ while ${\mbox{Tr}}(A_{i}X_{0})-b_{i}\tau_{0},i=1,\ldots,m$, and $\kappa_{0}-b^{T}y_{0}$ are generally nonzero.
Algorithm 4.1 can be written in an equivalent way as Algorithm 4.2, which we present below. We use Algorithm 4.2 to solve the sSDLCP. Before describing the algorithm, we first define an analogous (narrow) neighborhood of the central path of the SDLCP representation:
$\mathcal{N}_{1}(\beta,{{\color[rgb]{0,0,0}\hat{\mu}}}):=\{({{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}})\in S^{n_{1}}_{++}\times S^{n_{1}}_{++}\ ;\ \|({{\color[rgb]{0,0,0}\hat{Y}}})^{1/2}{{\color[rgb]{0,0,0}\hat{X}}}({{\color[rgb]{0,0,0}\hat{Y}}})^{1/2}-{{\color[rgb]{0,0,0}\hat{\mu}}}I\|_{F}\leq\beta{{\color[rgb]{0,0,0}\hat{\mu}}},{{\color[rgb]{0,0,0}\hat{\mu}}}={\rm{Tr}}({{\color[rgb]{0,0,0}\hat{X}}}{{\color[rgb]{0,0,0}\hat{Y}}})/n_{1}\}.$
We again have a system of equations for $(\Delta{{\color[rgb]{0,0,0}\hat{X}}},\Delta{{\color[rgb]{0,0,0}\hat{Y}}})\in S^{n_{1}}\times S^{n_{1}}$ that plays an important role in the algorithm, just like the system of equations (86)-(90) for Algorithm 4.1:
$({{\color[rgb]{0,0,0}\hat{Y}}})^{1/2}({{\color[rgb]{0,0,0}\hat{X}}}\Delta{{\color[rgb]{0,0,0}\hat{Y}}}+\Delta{{\color[rgb]{0,0,0}\hat{X}}}{{\color[rgb]{0,0,0}\hat{Y}}})({{\color[rgb]{0,0,0}\hat{Y}}})^{-1/2}+({{\color[rgb]{0,0,0}\hat{Y}}})^{-1/2}(\Delta{{\color[rgb]{0,0,0}\hat{Y}}}{{\color[rgb]{0,0,0}\hat{X}}}+{{\color[rgb]{0,0,0}\hat{Y}}}\Delta{{\color[rgb]{0,0,0}\hat{X}}})({{\color[rgb]{0,0,0}\hat{Y}}})^{1/2}$
$=2(\sigma{{\color[rgb]{0,0,0}\hat{\mu}}}I-({{\color[rgb]{0,0,0}\hat{Y}}})^{1/2}{{\color[rgb]{0,0,0}\hat{X}}}({{\color[rgb]{0,0,0}\hat{Y}}})^{1/2}),$ (104)
${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}(\Delta{{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}(\Delta{{\color[rgb]{0,0,0}\hat{Y}}})=-\overline{r}.$ (105)
Note that in Algorithm 4.2, it can be shown that $(\Delta{{\color[rgb]{0,0,0}\hat{X}}},\Delta{{\color[rgb]{0,0,0}\hat{Y}}})$ obtained by solving (104), (105) always exists and is unique [23, 26].
@@
(See Algorithm 4.1 of [26]; Algorithm 2.1 of [23]) Given $\epsilon>0$, and $\beta_{1}<\beta_{2}$ with $\beta_{2}^{2}/(2(1-\beta_{2})^{2})\leq\beta_{1}<\beta_{2}<\beta_{2}/(1-\beta_{2})<1$. Choose $({{\color[rgb]{0,0,0}\hat{X}}}_{0},{{\color[rgb]{0,0,0}\hat{Y}}}_{0})\in S^{n_{1}}_{++}\times S^{n_{1}}_{++}$ such that
${{\color[rgb]{0,0,0}\hat{X}}}_{0}=\left[\begin{array}[]{cc}X_{0}&0\\ 0&\tau_{0}\end{array}\right],\quad{{\color[rgb]{0,0,0}\hat{Y}}}_{0}=\left[\begin{array}[]{cc}Y_{0}&0\\ 0&\kappa_{0}\end{array}\right],$ (110)
@@
Set ${{\color[rgb]{0,0,0}\hat{X}}}={{\color[rgb]{0,0,0}\hat{X}}}_{k}$, ${{\color[rgb]{0,0,0}\hat{Y}}}={{\color[rgb]{0,0,0}\hat{Y}}}_{k}$, and define
$r:={{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}}).$
@@
If $\max\{{\rm{Tr}}({{\color[rgb]{0,0,0}\hat{X}}}{{\color[rgb]{0,0,0}\hat{Y}}})/({{\color[rgb]{0,0,0}\hat{X}}})^{2}_{n_{1},n_{1}},\|r/{{\color[rgb]{0,0,0}\hat{X}}}_{n_{1},n_{1}}\|\}\leq\epsilon$, then terminate with the solution $({{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}})$. If ${{\color[rgb]{0,0,0}\hat{X}}}_{n_{1},n_{1}}$ is sufficiently small, terminate with no optimal solutions to the sLSDFP.
@@
[Predictor Step] Find the solution $(\Delta{{\color[rgb]{0,0,0}{\hat{X}}}}_{p},\Delta{{\color[rgb]{0,0,0}\hat{Y}}}_{p})$ of the linear system (104), (105), with $\sigma=0$, $\overline{r}=r$.
@@
${{\color[rgb]{0,0,0}\overline{\hat{X}}}}={{\color[rgb]{0,0,0}\hat{X}}}+\overline{\alpha}\Delta{{\color[rgb]{0,0,0}\hat{X}}}_{p},\ \ {{\color[rgb]{0,0,0}\overline{\hat{Y}}}}={{\color[rgb]{0,0,0}\hat{Y}}}+\overline{\alpha}\Delta{{\color[rgb]{0,0,0}\hat{Y}}}_{p},$
@@
$\delta$ $=$ $\frac{1}{{{\color[rgb]{0,0,0}\hat{\mu}}}}\|({{\color[rgb]{0,0,0}\hat{Y}}})^{1/2}\Delta X_{p}^{1}\Delta Y_{p}^{1}({{\color[rgb]{0,0,0}\hat{Y}}})^{-1/2}\|_{F},$ (113)
@@
$\alpha_{2}=\max\{\tilde{\alpha}\in[0,1]\ ;\ ({{\color[rgb]{0,0,0}\hat{X}}}+\alpha\Delta{{\color[rgb]{0,0,0}\hat{X}}}_{p},{{\color[rgb]{0,0,0}\hat{Y}}}+\alpha\Delta{{\color[rgb]{0,0,0}\hat{Y}}}_{p})\in\mathcal{N}_{1}(\beta_{2},(1-\alpha){{\color[rgb]{0,0,0}\hat{\mu}}})\ \forall\ \alpha\in[0,\tilde{\alpha}]\}.$